OpenStudy (anonymous):

Solve the proportion 15t/5=2t+3/6

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Cross multiply... 15t*6=5(2t+3) 90t=10t+15 know what to do next?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got t=0.1875

OpenStudy (anonymous):

Thats right.

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

You're welcome =D